OpenStudy (anonymous):

Solve 15x^(4)-34x^(2)+15=0

OpenStudy (anonymous):

the solution is x=?

OpenStudy (campbell_st):

this is an equation that can be reduced to a quadratic let u = x^2 15u^2 - 34u + 15 use the general quadratic formula will give \[u = (34 \pm \sqrt{256}) / 30 \] u = 5/3 and 3/5 so then x^2 = 5/3 and 3/5 \[x = \pm \sqrt{5/3}, \pm \sqrt{3/5}\]